# Understanding Differential Transformer Unchains Pretrained Self-Attentions

### v2

@@ 1 Introduction
Building on these insights, we present Dex (Differential Extension), an efficient framework that injects the strengths of Diff Transformer into a pretrained LLM without training from scratch. Unlike most finetuning methods that fit the model to downstream data, Dex is an architectural adaptation strategy that introduces a key mechanism from a different architecture to a pretrained model, conceptually similar to MHA2MLA [37]. Specifically, Dex operates by reusing the pretrained softmax attention scores (softmax$(\mathbf{QK}^{T})$) and applying its learnable differential mechanism to the output value matrix (softmax$(\mathbf{QK}^{T})\mathbf{V}$), making the adaptation lightweight (in both training and inference) yet effective, as demonstrated empirically. To facilitate stable and performant transition, we introduce additional techniques for head selection and $\lambda$-annealing, which controls the critical balance between original knowledge and incoming architectural changes. We validate Dex across multiple model families (Llama-3 [25] and Qwen-2.5 [83]) and scales (0.5B-8B), using diverse benchmarks such as language modeling [24, 76, 78], key information retrieval [38] and in-context learning [6]. Dex consistently achieves significant gains using less than 0.01% the size of the original training data (<1B tokens), without incurring nontrivial test-time overhead.
@@ 2 How Does Differential Transformer Work?
In this section, we systematically analyze the internal mechanics of Diff Transformer. Since the original weights are not publicly available at the time of writing, we train a Diff Transformer on a similar data mix to carry out our analyses. Please refer to Appendix F.2 for full details.
@@ 3 Differential Extension
def MHA(X,W_q,W_k,W_v,f_D,W_o,$\lambda$,hs):
O = [Attn(X,...,$\lambda$,do=(i in hs)) for i in range(h)] # hs: selected heads
@@ 4.1 Language Modeling Evaluation
We apply Dex to Llama-3.1-8B [25], Llama-3.2-3B/1B [55], and Qwen-2.5-1.5B/0.5B [83]. As the original pretraining data for these models is unavailable, we build a custom corpus of web pages, papers, encyclopedias, and code from open datasets [43, 81], similar to OLMo [61]. This corpus contains 887M tokens (Llama-3 tokenizer), less than 0.01% of the models’ original pretraining data size. Although Dex is not presented as a fine-tuning method, we compare against baselines trained on the same data—including parameter-efficient tuning (PEFT; LoRA [31], PiSSA [53]) and full fine-tuning (FT; Galore [89], APOLLO [92])—to control for the influence of this corpus. Direct comparison with original Diff Transformer is limited by unavailable pretrained weights, which we defer to Appendix along with other details. For head selection we simply set $k$ to be half the total number of heads for each model, and we adopt the $\lambda_{init}$ from [84] (we provide ablations in Appendix B.3). For PiSSA we report $r=32$ case as this yields good results.
@@ Results
We report performances on 11 widely used language modeling benchmarks [15, 77, 75, 85, 57, 8, 67] using [24]. As shown in Table 1, Dex achieves significant improvements across model sizes and families. Given the discrepancy between our training corpus, original pretraining data and the downstream tasks, it is natural to observe degradation after additional training in some cases. Nevertheless, Dex demonstrates robust performance gains on the majority of benchmarks, even when other methods—all trained on the same corpus—exhibit performance drops. In particular, we attribute Dex’s strong performance on WSC to its enhanced anaphora resolution granted by the capacity to model negative relevance for incorrect antecedents, which aligns with our intuitions. Notably, although Dex only updates self-attentions, it consistently outperforms both PEFT and full fine-tuning even when full tuning steadily outperforms PEFT (e.g., Llama-1B). In Appendix C, we also show Dex’s effectiveness when applied to instruction tuning using Llama-8B model.
@@ 4.2 Key Information Retrieval
Increased attention to answer ratio in Table 2 further demonstrates that Dex effectively transfers the core capability of Diff attention: attention noise canceling. Despite being implicitly applied to the output value matrix, Dex notably alters the effective attention pattern to focus on relevant information. This result empirically supports our design choice, placing Dex on the sweetspot between efficiency and efficacy. We provide details in Appendix F.3.
@@
Figure 10: Multi-needle retrieval results. $N$: total number of needles, $R$: number of queries.
@@ 4.4 Ablation and Analysis
From Table 4, it is evident that incorporating entropy-based head selection combined with both learnable and annealed lambda methods yields the best performance, achieving the overall accuracy
Table 4: Ablation with head selection and lambda control strategies. ‘imp.’ refers to importance-based and ‘ent.’ stands for entropy-based.
@@
of 64.2%. Removing either component from lambda leads to noticeable performance drops, indicating the necessity of both. Additionally, both head selection strategies outperform the configuration without head selection, with the entropy-based strategy pushing the boundary further. The fact that choosing low entropy heads ($\downarrow$) underperforms further supports our design. These findings underline the complementary roles of the head selection and lambda annealing mechanisms in maximizing the effectiveness of Dex.
@@
Figure 11: Analysis on Dex. (a) Cosine similarity ($\mathrm{cosine}(\mathbf{O},\Delta)$) vs. Relative Norm ($||\Delta||/||\mathbf{O}||$, where $\Delta=\lambda f_{D}(\mathbf{O})$) (b) CKA of attention head output features (brighter means higher redundancy). (c) Training dynamics of learnable $\lambda$ under different initialization/annealing schemes.
Figure 12: Throughput comparison.
We verify the test-time efficiency of Dex by comparing the throughput (tokens per second) with base Llama and Diff Transformer (3B). In Fig.12, while Diff Transformer exhibits larger inefficiency with increased context length due to its compute-heavy attention operation, Dex remains competitive to original Llama.
@@
Finally, we evaluate the effect of training data sizes (Fig. 13). Notable gains appear with just 400M tokens, highlighting the lightweight nature of Dex. Since our training data lacks direct correlation with the downstream benchmarks, modest amount of data (<1B) is sufficient to elicit the full potential of Dex.
@@ 5 Conclusion
In this work, we study Diff Transformer. Based on the insights, we propose Dex, an effective architectural adaptation method to integrate the strengths of Diff attention into pretrained LLMs.
@@
For Diff Transformer analysis, we followed the original setup as much as possible, but different behaviors can emerge under different model scale, data composition, etc. Similarly, though Dex works well across model sizes, it has not been tested beyond 8B parameter scale.
@@ B.5 Qualitative Results
Appendix C Application to Instruction Tuning
To move beyond base pretrained models, we investigate whether Differential Extension (Dex) can likewise enhance performance on instruction-following tasks that are widely relevant to end-users. Instruction-following models typically undergo a much more elaborate training pipeline—e.g., supervised fine-tuning (SFT) followed by multiple rounds of reinforcement learning—and are usually trained on proprietary, high-quality datasets that differ substantially from publicly available corpora. Consequently, the experimental protocol established for base pretrained models cannot be directly transferred to the instruction-following scenario without modification.
To fairly assess the effect of Dex in this setting, we adopt two complementary strategies. First is integrating Dex on an instruction-tuned model, where we apply Dex to a publicly available checkpoint trained on an open-source instruction corpus OpenHermes-2.5 [70] [^2] https://huggingface.co/artificialguybr/Meta-Llama-3.1-8B-openhermes-2.5 to fully exclude the confounding effect of training data. We verify whether Dex can effectively be integrated into existing instruction-tuned models to further boost performances. Second is directly applying Dex to a base pretrained model as an instruction-tuning method, similarly using OpenHermes-2.5 as the training data but in a single stage without separate SFT. We examine if Dex itself can effectively induce instruction-following capabilities without relying on a prior instruction fine-tuning step. These two variants allow us to comprehensively inspect the efficacy of Dex in instruction tuning. Note that we include FT (further fine-tuning the open-source checkpoint on the same OpenHermes data for more iterations) as an additional baseline as we suspect the public checkpoint is underfitted, which helps us clearly disentangle the contributions of Dex from the general effects of further training.
Table 7: Instruction-tuning scores on 6 benchmarks. The first four rows correspond to the first approach (starting from an instruction-tuned checkpoint), while the bottom two rows correspond to the second (directly instruction-tuning the base pretrained model).
Model MMLU Arc-C IFEval MBPP++ GSM8K AGIEval AVG $\Delta$
Instruction-tuned
Base 62.9 78.3 46.8 68.3 71.1 32.2 59.9 -
+ LoRA 63.1 79.5 45.7 65.3 70.3 40.6 60.7 +0.8
+ FT 63.0 78.6 49.2 63.2 68.8 42.3 60.8 +0.9
+ Dex 63.1 77.7 57.2 64.8 74.3 40.7 63.0 +3.1
Pretrained
+ LoRA 63.7 70.5 42.0 65.3 57.4 35.4 55.7 -4.2
+ Dex 63.6 77.3 51.0 66.1 68.4 37.8 60.7 +0.8
Table 7 reports results on six representative task groups—language understanding [28], commonsense reasoning [15], instruction following [91], math [16], code generation [4], and general human task [90]. We observe that Dex delivers its largest gains on two of the benchmarks—IFEval and GSM8K. IFEval evaluates how well a model follows diverse instructions, and GSM8K tests mathematical problem-solving capabilities. Both tasks appear to benefit disproportionately from the differential mechanism of Dex, which concentrate its attention on the most relevant parts of the prompt, thus achieving the observed performance jump. AGIEval, on the other hand, improves substantially across all baselines (LoRA, FT, and Dex). We attribute this primarily to additional training rather than the architectural modifications themselves. Furthermore, when Dex is directly applied to a base pretrained model, it achieves significantly higher performance than LoRA. It also demonstrates performance comparable to more thoroughly instruction-tuned baselines (i.e., the SFT model with further fine-tuning (FT) or LoRA) from our first experimental strategy, even without Dex itself requiring a prior SFT stage. This indicates that Dex effectively and efficiently induces instruction-following capabilities.
Appendix D Comparison to Diff Transformer
A direct comparison with the original Diff Transformer model is not possible due to unavailable weights. Therefore, to establish a point of reference, we compare Dex with a Diff Transformer model that we trained ourselves at a smaller scale, following the procedures detailed in Appendix F.2. To set up this comparison, we first train two models from scratch on the exact same training data: (1) a standard transformer baseline using the Llama architecture (Llama), and (2) Diff Transformer model (Diff). Subsequently, we apply Dex to the Llama baseline using a small subset of the pretraining data (<1B tokens) to create the third model, simply noted Dex. We additionally train a separate Llama model from scratch with Dex attached from the beginning, to understand Dex’s architectural capacity beyond its original purpose of adaptation, which we refer to as Dex-S. We report the performance of these four models (Llama, Diff, Dex, Dex-S) on the 11 language modeling benchmarks in Table 8.
@@ Appendix D Comparison to Diff Transformer
Appendix E Efficiency Analysis
To evaluate inference efficiency, we benchmark throughput (tokens per second) for 3B-parameter versions of Llama, Diff Transformer, and Dex, presenting the results in Fig.12. Context lengths were varied from 1k to 64k tokens to cover a comprehensive range of use cases. All tests were conducted on a single NVIDIA A100-80GB GPU, utilizing PyTorch’s standard scaled dot-product attention implementation [^3] https://docs.pytorch.org/docs/stable/generated/torch.nn.functional.scaled_dot_product_attention.html. The reported throughputs are averaged over 30 batches, following an initial 5 warm-up batches.
Appendix F Implementation Details
@@
F.1 Language Modeling Evaluation
@@
F.2 Training Diff Transformer
@@
F.3 Approximating Effective Attention Scores for Dex Interpretability
@@
As with the pseudoinverse method, we aim to find an effective attention score matrix $\mathbf{X}$ such that $\mathbf{X}\mathbf{V}$ approximates the Dex output $\mathbf{O}^{\prime}=\mathbf{A}\mathbf{V}(I-\mathbf{W}_{D})$. Rather than a closed-form pseudoinverse solution, this approach directly optimizes for $\mathbf{X}$ for each input sample for which $\mathbf{O}^{\prime}$ and $\mathbf{V}$ are computed. For each sample, $\mathbf{X}$ is typically initialized (e.g., as the original attention scores $\mathbf{A}$) and then updated for 100 iterations using gradient descent (learning rate $1\times 10^{-3}$) to minimize a reconstruction loss with the form $||\mathbf{X}\mathbf{V}-\mathbf{O}^{\prime}||_{2}^{2}$.
@@
Appendix G Broader Impact

### v3

@@ 1 Introduction
[…]
@@ 2 How Does Differential Transformer Work?
In this section, we systematically analyze the internal mechanics of Diff Transformer. Since the original weights are not publicly available at the time of writing, we train a Diff Transformer on a similar data mix to carry out our analyses. Please refer to Appendix E.2 for full details.
@@ 3 Differential Extension
[…]
@@ 4.1 Language Modeling Evaluation
We apply Dex to Llama-3.1-8B [26], Llama-3.2-3B/1B [56], and Qwen-2.5-1.5B/0.5B [84]. As the original pretraining data for these models is unavailable, we build a custom corpus of web pages, papers, encyclopedias, and code from open datasets [44, 82], similar to OLMo [62]. This corpus contains 887M tokens (Llama-3 tokenizer), less than 0.01% of the models’ original pretraining data size. Although Dex is not presented as a fine-tuning method, we compare against baselines trained on the same data—including parameter-efficient tuning (PEFT; LoRA [32], PiSSA [54]) and full fine-tuning (FT; Galore [90], APOLLO [93])—to control for the influence of this corpus. Direct comparison with original Diff Transformer is limited by unavailable pretrained weights, and we defer evaluations in smaller settings to Appendix along with other details. For head selection we simply set $k$ to be half the total number of heads for each model, and we adopt the $\lambda_{init}$ from [85] (we provide ablations in Appendix B.3). For PiSSA we report $r=32$ case as this yields good results.
@@ Results
We report performances on 11 widely used language modeling benchmarks [16, 78, 76, 86, 58, 8, 68] using [25]. As shown in Table 1, Dex achieves significant improvements across model sizes and families. Given the discrepancy between our training corpus, original pretraining data and the downstream tasks, it is natural to observe degradation after additional training in some cases. Nevertheless, Dex demonstrates robust performance gains on the majority of benchmarks, even when other methods—all trained on the same corpus—exhibit performance drops. In particular, we attribute Dex’s strong performance on WSC to its enhanced anaphora resolution granted by the capacity to model negative relevance for incorrect antecedents, which aligns with our intuitions. Notably, although Dex only updates self-attentions, it consistently outperforms both PEFT and full fine-tuning even when full tuning steadily outperforms PEFT (e.g., Llama-1B).
@@ 4.2 Key Information Retrieval
Increased attention to answer ratio in Table 2 further demonstrates that Dex effectively transfers the core capability of Diff attention: attention noise canceling. Despite being implicitly applied to the output value matrix, Dex notably alters the effective attention pattern to focus on relevant information. This result empirically supports our design choice, placing Dex on the sweetspot between efficiency and efficacy. We provide details in Appendix E.3.
@@
Figure 10: Multi-needle retrieval results. NN: total number of needles, $R$: number of queries.
@@ 4.3 In-Context Learning
4.4 Application to Instruction Tuning
We investigate whether Dex can likewise enhance performance on instruction-following tasks. To fairly assess the effect of Dex, we adopt two complementary settings. First, we apply Dex on a publicly available instruction-tuned checkpoint trained on an open-source instruction corpus OpenHermes-2.5 [71] [^1] https://huggingface.co/artificialguybr/Meta-Llama-3.1-8B-openhermes-2.5 using the same training data (OH-2.5). This continued instruction tuning setting eliminates the confounding effect of training data and lets us verify whether Dex improves the performance of an existing instruct model. Second, we directly apply Dex to a base pretrained model as an instruction-tuning method itself, similarly using OH-2.5 but in single stage. We examine if Dex can effectively induce instruction-following capabilities without prior end-to-end instruction tuning. Note that we include FT (further fine-tuning the open-source checkpoint on the same OpenHermes data for more steps) as an additional baseline to alleviate the concern for underfitting, which clearly distinguishes the contribution of Dex from the benefit of more training steps.
Table 4: Instruction-tuning results on 8 benchmarks. The top four rows correspond to the first setting, while the bottom two rows correspond to the second.
Model MMLU Arc-C IFEval MBPP++ GSM8K AGIEval HumanEval Math500 AVG $\Delta$
Instruction-tuned
Base 62.9 78.3 46.8 68.3 71.1 32.2 44.5 13.4 52.2 -
+ LoRA 63.1 79.5 45.7 65.3 70.3 40.6 47.0 4.0 51.9 -0.3
+ FT 63.0 78.6 49.2 63.2 68.8 42.3 36.6 20.0 53.7 +1.5
+ Dex 63.1 77.7 57.2 64.8 74.3 40.7 47.6 19.2 55.6 +3.4
Pretrained
+ LoRA 63.7 70.5 42.0 65.3 57.4 35.4 45.7 2.0 47.8 -4.4
+ Dex 63.6 77.3 51.0 66.1 68.4 37.9 50.7 16.2 53.9 +1.7
Table 4 reports results on eight representative benchmarks that span language understanding [29], commonsense reasoning [16], instruction following [92], math [17, 29], code generation [4, 14], and general human task [91]. We observe that Dex delivers favorable results on diverse settings, significantly outperforming baselines on benchmarks like GSM8K, HumanEval and IFEval. When directly applied to a base pretrained model, Dex achieves notably higher performance than LoRA, demonstrating comparable performance to more heavily tuned baselines (top 3 rows) without any end-to-end SFT. These results indicate Dex’s effectiveness in inducing and reinforcing instruction-following capabilities efficiently.
@@ 4.5 Ablation and Analysis
Table 5: Ablation with head selection and lambda control strategies. imp. refers to importance-based and ent. stands for entropy-based.
@@
From Table 5, it is evident that incorporating entropy-based head selection combined with both learnable and annealed lambda methods yields the best performance, achieving the overall accuracy of 64.2%. Removing either component from lambda leads to noticeable performance drops, indicating the necessity of both. Additionally, both head selection strategies outperform the configuration without head selection, with the entropy-based strategy pushing the boundary further. The fact that choosing low entropy heads ($\downarrow$) underperforms further supports our design. These findings underline the complementary roles of the head selection and lambda annealing mechanisms in maximizing the effectiveness of Dex.
@@
[…]
We verify the test-time efficiency of Dex by comparing the throughput (tokens per second) and latency with base Llama and Diff Transformer (3B). Fig.12 clearly shows that both in terms of throughput and latency, Dex demonstrates superior inference time efficiency. While Diff Transformer exhibits increasing inefficiency with longer context due to its compute-heavy attention operation, Dex remains competitive to the original Llama baseline thanks to its lightweight design.
(a) Throughput comparison.
(b) Latency Comparison.
Figure 12: Inference time efficiency analysis. We benchmark (a) throughput and (b) latency of three attention variants. While Diff attention costs significantly more compute at test-time compared to the original Llama attention, Dex incurs negligible overhead thanks to its lightweight design. All benchmarks are measured on a single Nvidia A100 GPU.
@@
Finally, we evaluate the effect of training data size on the application of Dex (Fig. 13). Notable gains appear with just 400M tokens, highlighting the lightweight nature of Dex. Since our training data lacks direct correlation with the downstream benchmarks, modest amount of data (<1B) is sufficient to elicit the full potential of Dex, and simply adding more general data yields diminishing returns in downstream tasks.
@@ 5 Conclusion
In this work, we study the internal mechanism of Diff Transformer to identify three key factors behind its empirical success: enhanced expressivity via negative attention, reduced redundancy among attention heads and improved optimization dynamics. Based on these insights, we propose Dex, an architectural adaptation method that efficiently integrates the empirical strengths of Diff Transformer into pretrained LLMs with standard transformer architecture. Diverse evaluation results confirm the effectiveness and versatility of Dex.
@@
For Diff Transformer analysis, we followed the original setup as much as possible, but different behaviors can emerge under different model scale, data composition, etc. Similarly, though Dex works well across model sizes, it has not been tested beyond 8B parameter scale. We leave it for future works.
Acknowledgement
N. Kwak was supported by NRF (2021R1A2C3006659) and IITP grants (RS-2021-II211343, RS-2022-II220320, RS-2025-25442338), all funded by the Korean Government.
@@ B.5 Qualitative Results
Appendix C Comparison to Diff Transformer
A direct comparison with the original Diff Transformer model is not possible due to unavailable weights. Therefore, to establish a point of reference, we compare Dex with a Diff Transformer model that we trained ourselves at a smaller scale, following the procedures detailed in Appendix E.2. To set up this comparison, we first train two models from scratch on the exact same training data: (1) a standard transformer baseline using the Llama architecture (Llama), and (2) Diff Transformer model (Diff). Subsequently, we apply Dex to the Llama baseline using a small subset of the pretraining data (<1B tokens) to create the third model, simply noted Dex. We additionally train a separate Llama model from scratch with Dex attached from the beginning, to understand Dex’s architectural capacity beyond its original purpose of adaptation, which we refer to as Dex-S. We report the performance of these four models (Llama, Diff, Dex, Dex-S) on the 11 language modeling benchmarks in Table 8.
@@ Appendix C Comparison to Diff Transformer
Appendix D Efficiency Analysis
To evaluate inference efficiency, we benchmark throughput (tokens per second) for 3B-parameter versions of Llama, Diff Transformer, and Dex, presenting the results in Fig.12(a). Context lengths were varied from 1k to 64k tokens to cover a comprehensive range of use cases. All tests were conducted on a single NVIDIA A100-80GB GPU, utilizing PyTorch’s standard scaled dot-product attention implementation [^3] https://docs.pytorch.org/docs/stable/generated/torch.nn.functional.scaled_dot_product_attention.html. The reported throughputs are averaged over 30 batches, following an initial 5 warm-up batches.
Appendix E Implementation Details
@@
E.1 Language Modeling Evaluation
@@
E.2 Training Diff Transformer
@@
E.3 Approximating Effective Attention Scores for Dex Interpretability
@@
As with the pseudoinverse method, we aim to find an effective attention score matrix $\mathbf{X}$ such that $\mathbf{X}\mathbf{V}$ approximates the Dex output $\mathbf{O}^{\prime}=\mathbf{A}\mathbf{V}(I-\mathbf{W}_{D})$. Rather than a closed-form pseudoinverse solution, this approach directly optimizes for $\mathbf{X}$ for each input sample for which $\mathbf{O}^{\prime}$ and $\mathbf{V}$ are computed. For each sample, $\mathbf{X}$ is typically initialized (e.g., as the original attention scores $\mathbf{A}$) and then updated for 100 iterations using gradient descent (learning rate $1\times 10^{-3}$) to minimize a reconstruction loss with the form $\mathbf{X}\mathbf{V}-\mathbf{O}^{\prime}||_{2}^{2}$.
@@
Appendix F Broader Impact